OpenStudy (anonymous):

solve the triangle A= 45, a= 31, b=27

OpenStudy (anonymous):



OpenStudy (anonymous):

use cos theorem to find AB

OpenStudy (anonymous):

31^2=(AB)^2+27^2-2(AB)*27*cos45

OpenStudy (anonymous):

AB=x x^2-2*27*cos45*x+27^2-31^2=0

OpenStudy (anonymous):

BTW what do you mean solve the triangle ?

OpenStudy (anonymous):

since x is length, x cannot be negative take positive one..

OpenStudy (anonymous):